Louisiana State University Health Sciences Center Administration & Finance
 

Thursday, June 14, 2007

Dental School : 

  • Zimmer continues the general cleaning/remediation on the 8th floor of the Clinic building and on the 2nd through 4th floors in the Administration building. 
    • The 4th and 5th floor viable sample test results from the air handlers came back today. 
      • Both units on the 4th floor (4A & 4B) passed. 
      • Unit E-5B on the 5th floor also passed.  
      • Unit E-5A failed and the coils need to be re-cleaned.
      • The sampling results allowed both air handling units on the 4th floor and one of the two units on the 5th floor in the Clinic building to be re-started.  
    • Area air sample tests have been successfully passed on the 1st through 7th floors in the Clinic building. 
    • Sampling results for the air handling units on the 6th and 7th floors of the Clinic building are expected early next week. 
    • Zimmer was asked to place some negative air machines on the 4th floor and run flexible tube ducts up to the 6th and 7th floor to help cool those floors temporarily until we get the sample results back early next week. 
  • The demolition related to the Central Sterilization installation continued.
    • The main interior wall has been completely removed.
  • The new ceiling grid is being installed on the 3rd through 5th floors of the Clinic Building . 
    • Ceiling tiles are in place on the 2nd floor and partially installed on the 3rd floor. 
    • Flex duct and supply diffusers are being installed on the 2nd floor.
    • Frischhertz is re-installing lights on the 3rd and 4th floors.  They have started the repairs to the damaged junction boxes and conduit on the 5th floor.
    • Firetrol continues to reposition and add sprinkler heads as necessary.
  • The emergency generator was successfully tested and training on the equipment was conducted for 5 LSUHSC Facility Services employees.
  • Boe-Tel continues to run the new network and data lines.
  • General cleanup around the campus has begun with the removal of debris from the Power Plant.

1542 Tulane Ave.:

  • Arc Mechanical is installing one of the new air compressors in the rear of the basement.
  • Ordes Electrical is installing disconnects for the new air compressors and air driers.
  • Stratos Elevators continues installing the floor call buttons.  
    • The elevator inspection will have to be redone when the governor cable is replaced – the wrong one was sent and installed.

Landscaping Project:

  • Today all of the crepe myrtles were planted around both the Lions/LSUHSC Clinic Building and the CSRB. 
    • The final grading was completed around CSRB. 
    • Four additional sprinkler zones were activated.

MEB Cafeteria Remodeling:

  • All of the HVAC duct work has been fabricated and installed.
  • The circular terrazzo floor section has been poured in the center of the space.
  • Significant progress has been made in building the new stud walls and the fur downs in the ceiling.

Student Parking Lot Re-Surfacing:

  • Work begins tomorrow on a contract to pave Student Lot #2 and #3 with asphalt.